OpenStudy (anonymous):

Solve the equation. 5x - (-5x - 1) = -8 Select one: a. b. c. d.

OpenStudy (jmark):

5x - (-5x - 1) = -8 5x +5x+1 = -8 10x+1 = -8 10x = -8-1 10x = -9 x = -9/10 Get CBSE and ICSE help http://scoremore.beep.com/